the average grade of 30 students in their statistics and probability examination is 85 with a standard deviation of 5. find the interval for the true mean at 95% interval


1
Expert's answer
2022-05-13T05:14:51-0400

The critical value for "\alpha = 0.05" and "df = n-1 = 29" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.04523."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(85-2.04523\times\dfrac{5}{\sqrt{30}},85+2.04523\times\dfrac{5}{\sqrt{30}})"

"=(83.133, 86.867)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "83.133 < \mu < 86.867," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(83.133, 86.867)."